# Double-coset zeta functions for groups acting on trees

### v1

@@ 7.3. The behaviour at $s=-1$ and the Ihara zeta function of a weighted graph
where $I$ is the identity matrix of dimension $|E\Gamma|$ and $T=[T(a,b)]_{a,b\in E\Gamma}\in\mathrm{Mat}_{n}(\mathbb{R})$ is defined as $T(a,b)=W(a,b)$ for all $a,b\in E\Gamma$ (assuming to have set a total order on $E\Gamma$). The matrix $T$ is called the Bass operator of $(\Gamma,W)$ (cf. [9, Definition 3.10]). Note that (7.26) gives a meromorphic continuation of $Z_{(\Gamma,W)}(x)$ to $\mathbb{C}$.
@@
Let $(\Gamma,\omega)$ be an edge-weighted graph satisfying Setting [$\Gamma$]. Let $\Gamma_{1},\Gamma_{2}$ be subgraphs of $\Gamma$ satisfying $\Gamma=\Gamma_{1}\cup\Gamma_{2}$ and such that $\Gamma_{1}\cap\Gamma_{2}$ is a $1$-segment with edge set $\{a,\bar{a}\}$. Assume also that $t(a)$ and $o(a)$ are terminal vertices in $\Gamma_{1}$ and $\Gamma_{2}$, respectively. Let $W$, $W_{1}$ and $W_{2}$ be the transition weights defined in Example 7.11 on $\Gamma$, $\Gamma_{1}$ and $\Gamma_{2}$, respectively. Then,
@@
Denote by $\mathcal{E}(-1)$, $\mathcal{E}_{1}(-1)$, $\mathcal{E}_{2}(-1)$ the Bass operators of $\Gamma$, $\Gamma_{1}$ and $\Gamma_{2}$ at $-1$. Let also $I,I_{1}$ and $I_{2}$ denote the identity matrices with complex entries of dimension $|E\Gamma|$, $|E\Gamma_{1}|$ and $|E\Gamma_{2}|$, respectively. By Theorem 5.12 and Example 7.11,
(7.27) $\mathcal{Z}_{\Gamma,a\to a}(-1)=Z_{(\Gamma,W)}(1)\cdot\det(I-M).$
@@
(7.33) $\begin{array}[]{c}M=\begin{bmatrix}\tilde{M}_{1}&0\\ 0&\tilde{M}_{2}\end{bmatrix};\,M_{1}=\begin{bmatrix}\tilde{M}_{1}&0\\ 0&M(\bar{a},\bar{a})\end{bmatrix};\,M_{2}=\begin{bmatrix}M(a,a)&0\\ 0&\tilde{M}_{2}\end{bmatrix}.\end{array}$
@@
(7.34) $\begin{split}\det(I-M)&=\det(\tilde{I}_{1}-\tilde{M}_{1})\cdot\det(\tilde{I}_{2}-\tilde{M}_{2})\\ &=\frac{\det(I_{1}-M_{1})\cdot\det(I_{2}-M_{2})}{(1-M(\bar{a},\bar{a})(1-M(a,a)))}.\end{split}$

### v2

@@ 7.3. The behaviour at s=1s=-1 and the Ihara zeta function of a weighted graph
where $I$ is the identity matrix of dimension $|E\Gamma|$ and $T=[T(a,b)]_{a,b\in E\Gamma}\in\mathrm{Mat}_{n}(\mathbb{R})$ is defined as $T(a,b)=W(a,b)$ for all $a,b\in E\Gamma$ (assuming to have set a total order on $E\Gamma$). The matrix $T$ is called the Bass operator of $(\Gamma,W)$ (cf. [11, Definition 3.10]). Note that (7.26) gives a meromorphic continuation of $Z_{(\Gamma,W)}(x)$ to $\mathbb{C}$.
@@
Let $(\Gamma,\omega)$ be an edge-weighted graph satisfying Setting $\Gamma$]. Let $\Gamma_{1},\Gamma_{2}$ be subgraphs of $\Gamma$ satisfying $\Gamma=\Gamma_{1}\cup\Gamma_{2}$ and such that $\Gamma_{1}\cap\Gamma_{2}$ is a $1$-segment with edge set $\{a,\bar{a}\}$. Assume also that $t(a)$ and $o(a)$ are leaves in $\Gamma_{1}$ and $\Gamma_{2}$, respectively. Let $W$, $W_{1}$ and $W_{2}$ be the transition weights defined in Example 7.11 on $\Gamma$, $\Gamma_{1}$ and $\Gamma_{2}$, respectively. Then,
@@
Denote by $\mathcal{E}(-1)$, $\mathcal{E}_{1}(-1)$, $\mathcal{E}_{2}(-1)$ the Bass operators of $\Gamma$, $\Gamma_{1}$ and $\Gamma_{2}$ at 1-1, respectively. Let also $I,I_{1}$ and $I_{2}$ denote the identity matrices with complex entries of dimension $|E\Gamma|$, $|E\Gamma_{1}|$ and $|E\Gamma_{2}|$, respectively. By Theorem 5.14 and Example 7.11,
(7.27) $\mathcal{Z}_{\Gamma,a\to a}(-1)=Z_{(\Gamma,W)}(1)\cdot\det(I-M),$
@@
[…]
@@
(7.34) $\begin{split}\det(I-M)&=\det(\tilde{I}_{1}-\tilde{M}_{1})\cdot\det(\tilde{I}_{2}-\tilde{M}_{2})\\ &=\frac{\det(I_{1}-M_{1})\cdot\det(I_{2}-M_{2})}{(1-M(\bar{a},\bar{a}))(1-M(a,a))}.\end{split}$
@@
Declaration of AI-assisted technologies in the writing process
During the preparation of this work, the author used “Writefull” to check spelling and grammar. After using this tool, the author reviewed and edited the content as needed and takes full responsibility for the content of the published article.